OpenStudy (anonymous):

1. What are the perimeter and area of a rectangle with a length of 20 in. and a width of 15 in.? (1 point) 1. P = 35 in., A = 300 in.2 2. P = 70 in., A = 300 in.2 3. P = 300 in., A = 70 in.2 4. P = 35 in., A = 200 in.2

OpenStudy (anonymous):

Perimeter of rectangle is 2*(length+breadth) i.e.2*(20+15) or P=2*35=70 in. Area of rectangle is length*breadth A=20*15 A=300 in.2 option a)